If (P − 1) is directly proportional to (Q + 1)^2
and P = 19 when Q = 2, find Q when P = 51.

Answer by Boreal(15235)

(p-1)=k((Q+1)^2
(19-1)=k(2+1)^2
18=9k
k=2
-
(51-1)=2(Q+1)^2
25=(Q+1)^2
Q=4